OpenStudy (anonymous):

17. What are the solutions of the equation? z² – 6z – 27 = 0 (1 point) A) 3, 9 B) 3, –9 C) –3, 9 D) –3, –9 18. Tasha is planning an expansion of a square flower garden in a city park. If each side of the original garden is increased by 7 m, the new total area of the garden will be 144 m². Find the length of each side of the original garden. (1 point) A) 19 m B) 12 m C) 5 m D) √5 m

OpenStudy (anonymous):

For the first one... the answer is B -9 x 3 = -27 -9+3 = -6

OpenStudy (anonymous):

the second answer is B

OpenStudy (anonymous):

sqrt of 144 is 12 which is the side of the square after being increased by 7... so 12 - 7 = the old length of the side which is 5

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

so for the second one...the answer is C?

OpenStudy (anonymous):

yupp :D

OpenStudy (anonymous):

thanks...can i message you a few more questions?

OpenStudy (anonymous):

lol i'm actually getting off os... sorry booboo

OpenStudy (anonymous):

oh ok :(

OpenStudy (anonymous):

:( sowwy :(

OpenStudy (anonymous):

its fine